# Distributionally Robust Markov Games with Average Reward

### v1

@@ 3 Hardness in Average Reward Nash Equilibrium Identification
Additional hidden environment player. The first major difference between a vanilla non-robust MG and a DR-MG is the introduction of an additional player, which selects the transition kernel or the underlying game model to diminish the gains of the other players Zhang et al. (2020). This additional player fundamentally changes the structure of non-robust MGs. For example, in the extensively studied special case of two-player zero-sum MGs, where $r_{1}=-r_{2}=r$, one fundamental structure of standard MGs is the Min-Max duality:
$\max_{\pi_{1}}\min_{\pi_{2}}V^{(\pi_{1},\pi_{2})}_{\mathsf{P},r}=\min_{\pi_{2}}\max_{\pi_{1}}V^{(\pi_{1},\pi_{2})}_{\mathsf{P},r}$, (3)
which directly implies the existence of a Nash Equilibrium (NE) Shapley (1953). Such results are further extended to non-robust average-reward MG studies through the limit method, which sets the discount factor $\gamma\to 1$ Filar and Vrieze (2012); Sennott (1994); Rieder (1995).
@@
$\max_{\pi_{1}}\min_{\pi_{2}}\min_{\mathsf{P}\in\mathcal{P}}V^{(\pi_{1},\pi_{2})}_{\mathsf{P},r}\neq\min_{\pi_{2}}\max_{\pi_{1}}\max_{\mathsf{P}\in\mathcal{P}}V^{(\pi_{1},\pi_{2})}_{\mathsf{P},r}$. (4)
@@
Consequently, even through the uniform convergence of the discounted robust value functions enables the interchanging of the order of $\max,\min,\lim$ Wang et al. (2023a), it does not imply the existence of NE.
Complicated structure of the average reward. The second major challenge stems from the inherent complexity of the robust average reward. When compared to the Bellman equations for the robust discounted reward Iyengar (2005); Nilim and El Ghaoui (2004); Xu and Mannor (2010) and the non-robust average reward, the robust average reward has a more complicated Bellman equation Wang et al. (2023a); Wang, Zou, and Wang (2024). Primarily, its solvability is not yet justified (an incorrect proof is given in Wang et al. (2023a)), therefore making the connection between optimal policies and the robust Bellman equation unclear. Moreover, even if it is solvable, its solution is not unique, even up to constant vectors Wang et al. (2023b). Whereas, in the non-robust setting, the average-reward Bellman equation has a unique solution up to constant vectors Puterman (1994). Therefore, despite the existence of the NE being proved for both discounted robust MGs Kardeş, Ordóñez, and Hall (2011); Zhang et al. (2020) and average-reward non-robust MGs Guo and Yang (2008); Hernández-Lerma and Lasserre (2000); Guo and Hernández-Lerma (2003); Zheng and Guo (2024); Nowak (1999); Iwase, Tanaka, and Wakuta (1976); Küenle and Schurath (2003); Tanaka and Wakuta (1977); Jaśkiewicz and Nowak (2001), the existence of the NE for average-reward robust MGs remains to be seen.
By identifying these challenges, we aim to develop a novel method to identify the Nash Equilibrium of average-reward distributionally robust MGs.
4 Solvability of the Robust Bellman Equation
In this section, we develop a deeper study of the single-agent robust Bellman equation, which will serve as an essential role in our following studies of distributionally robust games. Despite the extensive study of the robust Bellman equation, e.g., Wang et al. (2023a, b); Xu, Mondal, and Aggarwal (2025); Xu, Ganesh, and Aggarwal (2025); Roch et al. (2025); Roch, Atia, and Wang (2025); Chen, Wang, and Si (2025), whether this equation is solvable is in question. Hence, in this section, we aim to close this gap by examining the solvability of the robust Bellman equation.
To facilitate our analysis, we adopt the following standard irreducible assumption from (robust) average-reward studies Filar and Vrieze (2012); Xu, Mondal, and Aggarwal (2025); Xu, Ganesh, and Aggarwal (2025). Under it, the average reward is a constant and is independent from the initial state.
Assumption 4.1.
@@ 4 Solvability of the Robust Bellman Equation
For any deterministic policy $\pi$ and any transition kernel $\mathsf{P}\in\mathcal{P}$, the induced Markov chain $(\mathcal{S},\mathcal{A},\mathsf{P}^{\pi})$ is irreducible, i.e., all states belong to a single recurrent class [^2] We naturally extend this assumption to the multi-agent setting: $\mathsf{P}^{\pi}$ is irreducible for any joint deterministic policy and any $\mathsf{P}\in\mathcal{P}$.. Moreover, $\mathcal{P}$ is compact and convex.
Remark 4.2.
For convenience and to simplify the problem, we adopt the irreducibility assumption, which is slightly stronger than the standard unichain assumption Wang et al. (2023a, b). However, our results can be extended to the unichiain cases (see Section F in the Appendix).
We then derive the solvability of the robust Bellman equation under Assumption 4.1 as follows.
Theorem 4.3.
(1). (Robust Bellman Equation)
Consider an arbitrary uncertainty set, $\mathcal{P}$, satisfying Assumption 4.1, and denote any worst-case transition kernel w.r.t. $g^{\pi}_{\mathsf{P}}$ by $\mathsf{P}_{w}\triangleq\arg\min_{\mathsf{P}\in\mathcal{P}}g^{\pi}_{\mathsf{P}}$. Then the gain/bias of $(\pi,\mathsf{P}_{w}):$ $(g^{\pi}_{\mathsf{P}_{w}},h^{\pi}_{\mathsf{P}_{w}})$ is a solution to the equation:
$V(s)=\sum_{a}\pi(a|s)(r(s,a)-g+\sigma_{\mathcal{P}^{a}_{s}}(V)),$ (5)
where $\sigma_{\mathcal{P}^{a}_{s}}(V)\triangleq\min_{P\in\mathcal{P}^{a}_{s}}PV$.
(2). (Optimal Robust Bellman Equation) The optimal robust Bellman equation also has a solution:
$V(s)=\max_{a}\{r(s,a)-g(s)+\sigma_{\mathcal{P}^{a}_{s}}(V)\}$ (6)
Our results therefore prove the solvability of the robust Bellman equation, solidifying the connection between the robust average reward and its solutions. We note that the robust Bellman equation can have a more complicated structure than the non-robust one: any worst-case kernel $\mathsf{P}_{w}$ will result in a solution, hence (5) can have non-unique solutions even up to constant vectors. Our proof techniques are also different from the non-robust ones, which rely on the Laurent series of the non-robust average reward Puterman (1990), whereas we directly show the constructed pair $(g^{\pi}_{\mathsf{P}_{w}},h^{\pi}_{\mathsf{P}_{w}})$ is a solution.
These results and connections are of fundamental importance. As we will later show, solving the single-agent Bellman equations will be an essential step of showing the existence of a Nash Equilibrium.
5 Average-Reward Nash Equilibrium
In this section, we study the existence of a Nash Equilibrium in a distributionally robust Markov game under the average-reward criterion. As discussed in Section 3, neither the limit method nor the Min-Max duality is applicable in average-reward DR-MGs, hence we propose a direct approach by studying the robust best response policy.
Definition 5.1.
@@ 5 Average-Reward Nash Equilibrium
That is, the best response policy is agent $i$’s optimal policy when other agents fix their policies. This notion and the following results are standard and extensively studied in game theory literature Osborne and Rubinstein (2004); Fudenberg and Tirole (1991):
@@
$\pi^{*}$ is a Nash Equilibrium if $\pi^{*}_{i}\in\text{BR}_{i}(\pi^{*}_{-i})$, $\forall i\in\mathcal{N}$; or equivalently, if $\pi^{*}$ is a fixed point of the BR mapping: $\pi^{*}\in\text{BR}(\pi^{*})$, where $\text{BR}(\pi)=\otimes_{i}\text{BR}_{i}(\pi^{*}_{i})$.
Lemma 5.3.
Kakutani (1941)(Kakutani’s Fixed Point Theorem) Let $X$ be a compact convex subset of $\mathbb{R}^{n}$ and let $f:X\to 2^{X}$ be a set-valued function for which, (1) for all $x\in X$, $f(x)$ is nonempty and convex; (2) $f$ is upper semi-continuous (i.e., for any converging sequences $\{x_{n}\}\to x$ and $\{y_{n}\}\to y$ such that $y_{n}\in f(x_{n})$ $\forall n$, it holds that $y\in f(x)$). Then there exists some $x^{*}\in X$ such that $x^{*}\in f(x^{*})$.
@@
If for any agent $i\in\mathcal{N}$ and any stationary joint policy $\pi_{-i}$, $\text{BR}_{i}(\pi_{-i})$ is nonempty, convex, and $\text{BR}_{i}(\cdot)$ is upper semi-continuous, then the BR also satisfies the conditions in Lemma 5.3.
@@ 5.1 Induced Distributionally Robust MDPs
Definition 5.5.
(Induced Distributionally Robust MDP) Given a distributionally robust Markov game $\mathcal{MG}=(\mathcal{N},\mathcal{S},\mathcal{A},\mathcal{P},r)$ and an agent $i\in\mathcal{N}$. The induced distributionally robust MDP of the $\mathcal{MG}$ by a joint policy $\pi_{-i}$ is defined as $\mathcal{M}(\mathcal{MG},i,\pi_{-i})\triangleq(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{-i}},r^{\pi_{-i}})$, where
@@
Clearly, $(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{-i}},r^{\pi_{-i}})$ is a single-agent distributionally robust MDP for agent $i$. The uncertainty set $\mathcal{P}^{\pi_{-i}}$ is $(s,a_{i})$-rectangular, due to the $(s,a)$-rectangularity of the underlying uncertainty set $\mathcal{P}$. Moreover, the irreducibility, convexity, and compactness are all preserved through inducing. Namely, if the distributionally robust Markov game $\mathcal{MG}$ satisfies the multi-agent extension of Assumption 4.1, then the induced distributionally robust MDP $\mathcal{M}(\mathcal{MG},i,\pi_{-i})$ satisfies Assumption 4.1. Thus all of the previous results for the single-agent setting are still valid under the induced distributionally robust MDP for a fixed policy $\pi_{-i}$, based on which we can show the existence of a NE.
@@ 5.2 Existence of a Nash Equilibrium
We first show the following convexity result.
@@
We further show the semi-continuity of the optimal policy set of the induced distributionally robust MDP.
@@
Note that in this result, the inducing policy $\pi^{n}_{-i}$ varies, thus resulting in a time-varying single-agent distributionally robust MDP. Hence, we need to reveal the inherent continuity of induced distributionally robust MDPs.
@@
(Existence of a Nash Equilibrium) For an average-reward distributionally robust Markov game satisfying Assumption 4.1, a stationary Nash equilibrium exists.

### v2

@@ 3 Hardness in Average Reward Nash Equilibrium Identification
Additional hidden environment player. The first major difference between a vanilla non-robust MG and a DR-MG is the introduction of an additional player, which selects the transition kernel or the underlying game model to diminish the gains of the other players (zhang2020robust). This additional player fundamentally changes the structure of non-robust MGs. For example, in the extensively studied special case of two-player zero-sum MGs, where $r_{1}=-r_{2}=r$, one fundamental structure of standard MGs is the Min-Max duality:
$\max_{\pi_{1}}\min_{\pi_{2}}V^{(\pi_{1},\pi_{2})}_{\mathsf{P},r}(\rho)=\min_{\pi_{2}}\max_{\pi_{1}}V^{(\pi_{1},\pi_{2})}_{\mathsf{P},r}(\rho),$ (3)
which directly implies the existence of a Nash Equilibrium (NE) (shapley1953stochastic). Such results are further extended to non-robust average-reward MG studies through the limit method, which sets the discount factor $\gamma\to 1$ (filar2012competitive; sennott1994zero; rieder1995average).
@@
$\max_{\pi_{1}}\min_{\pi_{2}}\min_{\mathsf{P}\in\mathcal{P}}V^{(\pi_{1},\pi_{2})}_{\mathsf{P},r}(\rho)\neq\min_{\pi_{2}}\max_{\pi_{1}}\max_{\mathsf{P}\in\mathcal{P}}V^{(\pi_{1},\pi_{2})}_{\mathsf{P},r}(\rho).$ (4)
@@
Consequently, even through the uniform convergence of the discounted robust value functions enables the interchanging of the order of $\max,\min,\lim$ (wang2023robust), it does not imply the existence of NE. And this approach cannot be applied in multi-player general-sum games.
Complicated structure of the average reward. The second major challenge stems from the inherent complexity of the (robust) average reward, which heavily depends on structures of the underlying MDPs. Firstly, even in the non-robust setting, Markov games with average reward do not necessarily admit a stationary NE (filar2012competitive; wikecek2015stationary; lozovanu2016stationary; thuijsman1997survey). And with model uncertainty, the robust average reward has a more complicated structure, including a non-linear robust Bellman equation (wang2023robust; wang2024modelfree) with non-unique solutions. The studies and understandings of robust average reward is also significantly limited compared to non-robust and discounted settings (wang2023robust; wang2023model; wang2025bellman). Therefore, despite the existence of the NE being proved for both discounted robust MGs (kardecs2011discounted; zhang2020robust) and average-reward non-robust MGs (guo2008new; hernandez2000zero; guo2003zero; zheng2024zero; nowak1999optimal; iwase1976markov; kuenle2003optimality; tanaka1977continuous; jaskiewicz2001optimality; wikecek2015stationary), these results and methods cannot be applied for the NE of average-reward DR-MGs.
To better illustrate the hardness of solving average reward DR-MGs, we first show the following impossibility result. The constructed DR-MG is essentially a multi-chain (puterman1994markov), indicating the necessity of the additional structural assumptions to solve average reward DR-MGs.
Lemma 3.1.
There exists a two-player DR-MG without any stationary NE under average reward.
4 NE of Irreducible DR-MGs
As discussed above, the existence of a NE cannot be guaranteed without any structural assumptions. In this section, we first study a relatively easier setting, where the underlying DR-MG is assumed to be irreducible.
Assumption 1.
@@ 4 NE of Irreducible DR-MGs
For any deterministic joint policy $\pi$ and any transition kernel $\mathsf{P}\in\mathcal{P}$, the induced Markov chain $(\mathcal{S},\mathcal{A},\mathsf{P}^{\pi})$ is irreducible, i.e., all states belong to a single recurrent class. Moreover, $\mathcal{P}$ is compact and convex.
Remark 1.
Irreducibility is a standard assumption (slightly stronger than the unichain assumption (wang2023robust; wang2023model; wang2024robust; roch2025finite; Roch2025reduction)) in robust average reward studies (xu2025finite; xu2025efficient). Under this assumption, the robust average reward $g^{\pi}_{\mathcal{P}}(\rho)=g^{\pi}_{\mathcal{P}}(s)$ is always a constant over all initial states, ensuring the tractability of robust average reward. Hence in this section, we omit the initial state and denote the robust average reward by $g^{\pi}_{\mathcal{P}}$.
As discussed in Section 3, neither the limit method nor the Min-Max duality is applicable in average-reward DR-MGs. In non-robust average reward MG studies, the NE can be proved by studying the induced stationary distribution (wikecek2015stationary). However, when considering robustness, the worst-case stationary distribution set becomes much more complicated. For example, it is non-convex (wang2022robust; zhang2024distributionally; ma2025rectified), whereas the convexity is necessary in non-robust proofs and hence the method cannot be adapted. We thus propose a direct approach by studying the best response policy.
Definition 1.
@@
That is, the best response policy is agent $i$’s optimal policy when other agents fix their policies. This notion and the following results are standard and extensively studied in game theory literature (osborne2004introduction; fudenberg1991game):
@@
$\pi^{\star}$ is a Nash Equilibrium if $\pi^{\star}_{i}\in\text{BR}_{i}(\pi^{\star}_{-i})$, $\forall i\in\mathcal{N}$; or equivalently, if $\pi^{\star}$ is a fixed point of the BR mapping: $\pi^{\star}\in\text{BR}(\pi^{\star})$, where $\text{BR}(\pi)=\otimes_{i}\text{BR}_{i}(\pi^{\star}_{-i})$.
Lemma 4.2 (Kakutani’s Fixed Point Theorem (kakutani1941generalization)).
Let $X$ be a compact convex subset of $\mathbb{R}^{n}$ and let $f:X\to 2^{X}$ be a set-valued function for which, (1) $f(x)$ is nonempty and convex $\forall x$; (2) $f$ is upper hemi-continuous (i.e., for any convergent sequences $\{x_{n}\}\to x$ and $\{y_{n}\}\to y$ such that $y_{n}\in f(x_{n})$, it holds that $y\in f(x)$). Then there exists some $x^{\star}\in X$ such that $x^{\star}\in f(x^{\star})$.
@@
If for any agent $i\in\mathcal{N}$ and any stationary joint policy $\pi_{-i}$, $\text{BR}_{i}(\pi_{-i})$ is nonempty, convex, and $\text{BR}_{i}(\cdot)$ is upper hemi-continuous, then the BR also satisfies the conditions in Lemma 4.2.
@@ 4.1 Induced Distributionally Robust MDPs
Definition 2 (Induced Distributionally Robust MDP).
Given a distributionally robust Markov game $\mathcal{MG}=(\mathcal{N},\mathcal{S},\mathcal{A},\mathcal{P},r)$ and an agent $i\in\mathcal{N}$. The induced distributionally robust MDP of the $\mathcal{MG}$ by a joint policy $\pi_{-i}$ is defined as $\mathsf{M}_{i}(\pi_{-i})\triangleq(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{-i}},r^{\pi_{-i}})$, where
@@
Clearly, $(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{-i}},r^{\pi_{-i}})$ is a single-agent distributionally robust MDP for agent $i$. The uncertainty set $\mathcal{P}^{\pi_{-i}}$ is $(s,a_{i})$-rectangular, due to the $(s,a)$-rectangularity of the underlying uncertainty set $\mathcal{P}$. Moreover, due to being induced, irreducibility, convexity, and compactness are all preserved. Namely, if the distributionally robust Markov game $\mathcal{MG}$ satisfies the multi-agent extension of Assumption 1, then the induced distributionally robust MDP $\mathsf{M}_{i}(\pi_{-i})$ satisfies Assumption 1. Thus all of the previous results for the single-agent setting are still valid under the induced distributionally robust MDP for a fixed policy $\pi_{-i}$, based on which we can show the existence of a NE.
However, even the single-agent average reward robust RL is not fully studied. Specifically, the most essential tool, its robust Bellman equation, is not well studied: its solvability is not yet justified (an incorrect proof is given in (wang2023robust)), therefore making the connection between optimal policies and the robust Bellman equation unclear. Although a huge body of existing work develops convergent algorithms, their convergence guaranties either rely on the wrong result in (wang2023robust), e.g., (wang2023model; Roch2025reduction), or bypass its solvability (xu2025finite; xu2025efficient; chatterjee2023solving; meggendorfer2025solving). We hence derive the solvability of the robust Bellman equation under Assumption 1 as follows.
Theorem 4.1.
(1). (Robust Bellman Equation). Under Assumption 1, denote any worst-case transition kernel w.r.t. $g^{\pi}_{\mathsf{P}}$ by $\mathsf{P}_{w}\triangleq\arg\min_{\mathsf{P}\in\mathcal{P}}g^{\pi}_{\mathsf{P}}$. Then the gain/bias of $(\pi,\mathsf{P}_{w}):$ $(g^{\pi}_{\mathsf{P}_{w}},h^{\pi}_{\mathsf{P}_{w}})$ is the unique solution to the equation:
$V(s)=\sum_{a}\pi(a|s)(r(s,a)-g+\sigma_{\mathcal{P}^{a}_{s}}(V)),$ (7)
where $\sigma_{\mathcal{P}^{a}_{s}}(V)\triangleq\min_{P\in\mathcal{P}^{a}_{s}}PV$; Moreover, for any solution $(g,V)$ to (7), it holds that $g=g^{\pi}_{\mathcal{P}},V=h^{\pi}_{\mathsf{P}_{w}}+c\textbf{e}$, for some constant $c$ and $\textbf{e}=(1,...,1).$
(2). (Optimal Robust Bellman Equation). The optimal robust Bellman equation also has a solution:
$V(s)=\max_{a}\{r(s,a)-g+\sigma_{\mathcal{P}^{a}_{s}}(V)\}.$ (8)
Our results therefore prove the solvability of the robust Bellman equation, solidifying the connection between the robust average reward and its solutions. We note that the robust Bellman equation can have a more complicated structure than the non-robust one: any worst-case kernel $\mathsf{P}_{w}$ will result in a solution, hence (7) can have non-unique solutions even up to constant vectors. Our proof techniques are also different from the non-robust ones, which rely on the Laurent series of the non-robust average reward (puterman1990markov), whereas we directly show the constructed pair $(g^{\pi}_{\mathsf{P}_{w}},h^{\pi}_{\mathsf{P}_{w}})$ is a solution.
These results and connections are of fundamental importance in showing the existence of a Nash Equilibrium.
@@ 4.2 Existence of a Nash Equilibrium
With the results derived for single-agent robust Bellman equations, we can then show the following results.
@@
We further show the hemi-continuity of the optimal policy set of the induced distributionally robust MDP.
@@
Note that in this result, the inducing policy $\pi^{n}_{-i}$ varies, thus resulting in a distributionally robust MDP sequence with different dynamics. Hence, we need to reveal the inherent continuity of induced distributionally robust MDPs.
Remark 2.
The proofs of these results are based on the correspondence between the optimal policies and the solutions to the Bellman equations, heavily relying on Theorem˜4.1. In the next section when we study a weaker setting, such correspondence does not hold anymore and the proof derived cannot be applied there.
@@
(Existence of a Nash Equilibrium) For an average-reward distributionally robust Markov game satisfying Assumption 1, a stationary Nash equilibrium exists.
@@
5 NE of Weakly Communicating DR-MGs
In this section, we further extend our studies to a more general setting of weakly communicating DR-MGs.
Assumption 2 (Weakly Communicating DR-MG).
$\mathcal{P}$ is convex and compact. For any transition kernel $\mathsf{P}\in\mathcal{P}$, and for any player $i$, the state space can be partitioned into two distinct sets $\mathcal{S}=\mathcal{S}_{0}^{i}\cup\mathcal{S}_{1}^{i}$, such that (1). Any state in $\mathcal{S}_{0}^{i}$ is transient under every stationary joint policy; And (2). For any two states $s,s^{\prime}\in\mathcal{S}_{1}^{i}$, and any policy $\pi_{-i}$, there exists some policy $\pi_{i}$ and $N$ such that $\mathbb{P}(S_{N}=s^{\prime}|S_{0}=s,(\pi_{i},\pi_{-i}),\mathsf{P})>0$.
This assumption is adapted from the non-robust settings (wikecek2015stationary; bertsekas2011dynamic). The weakly communicating assumption is strictly weaker than ˜1, as it only requires accessibility under some policy, instead of every policy. Moreover, under this assumption, only the optimal robust average reward $g^{\star}_{\mathcal{P}}(s)$ is a constant over initial states, whereas $g^{\pi}_{\mathcal{P}}(s)$ is not.
The solvability of the optimal robust Bellman equation is derived recently in (wang2025bellman), as follows.
Lemma 5.1.
(wang2025bellman) Under ˜2, (8) has a solution $(g^{\star},u^{\star})$, and $g^{\star}=g^{\star}_{\mathcal{P}}(\rho)$.
The results are strictly weaker than the ones we derived for irreducible DR-MGs. Firstly, Theorem˜4.1 ensures the solvability of the robust Bellman equation with any policy, whereas Lemma˜5.1 only consider the optimal equation. Hence, a correspondence between the optimal robust policy and solutions of (8) cannot be developed as in irreducible settings, and the previous proofs cannot be applied to show the convexity and continuity of the best-response mapping.
To address this issue, we instead study another mapping specified through the optimal robust Bellman equation: for $s\in\mathcal{S}$, $\nu\in\Delta(\mathcal{A}_{i})$, and a joint policy $\pi_{-i}$, we define
$\Psi_{s}(\nu)\triangleq\sum_{a_{i}\in\mathcal{A}_{i}}\nu(a_{i})\big(r^{\pi_{-i}}_{i}(s,a_{i})-g^{\star}+\sigma_{(\mathcal{P}^{\pi_{-i}})^{a_{i}}_{s}}(u^{\star})\big),$
and set a map as the solution set of the Bellman equation:
$\mathcal{G}_{i}(\pi_{-i})\ \triangleq\ \bigl\{\nu:\Psi_{s}(\nu(\cdot|s))=\max_{\mu\in\Delta(\mathcal{A}_{i})}\Psi_{s}(\mu),\forall s\in\mathcal{S}\bigr\}.$
Based on the solvability result in Lemma˜5.1, we can show that $\mathcal{G}_{i}$ is nonempty. Different from irreducible cases, due to the existence of transient states, it is not guaranteed that an optimal policy can be obtained from a solution to the Bellman equations, hence $\mathcal{G}_{i}(\pi_{-i})$ may not contain all best-response policies. However, we can show that $\mathcal{G}_{i}\subseteq\text{BR}_{i}(\pi_{-i})$, which is sufficient to derive the existence of the NE. Further, if the map $\mathcal{G}(\pi)=\times_{i}\mathcal{G}_{i}(\pi_{-i})$ satisfies the conditions in Lemma˜4.2, it then admits a fixed point which is also a fixed point under the BR mapping, and hence implies the existence of NE. Our results are presented in the following theorem.
Theorem 5.1.
Under ˜2, the following statements hold. (1). $\mathcal{G}_{i}$ is nonempty, compact, and convex; (2). $\mathcal{G}_{i}$ is upper hemi-continuous; (3). $\mathcal{G}_{i}\subseteq\text{BR}_{i}(\pi_{-i})$; (4). A DR-MG always has a stationary NE.
The proofs are deferred to Appendix˜D, which rely on the strong duality of the robust average reward shown in (wang2025bellman), and are fundamentally different from the ones for Section˜4.